Consider four different stocks, all of which have a required return of 20 percent and a most recent dividend of $4.70 per shaPlease include work in the BA 2 Plus Calculator if possible with this question.
Answer #1

Dividend yield=D1/P0=D0*(1+g)/(D0*(1+g)/(r-g))

Capital gains yield=required return-dividend yield

1.
=4.70*1.1/(4.70*1.1/(20%-10%))=10.000%

2.
=4.70*1/(4.70*1/(20%-0%))=20.000%

3.
=4.70*0.95/(4.70*0.95/(20%+5%))=25.000%

4.
=4.70*1.2/(4.7*(1.2/1.2)+4.7*(1.2/1.2)^2+4.7*(1.2/1.2)^2*1.11/(20%-11%))=8.372%

5.
=20%-10.000%=10.00%

6.
=20%-20.000%=0.00%

7.
=20%-25.000%=-5.00%

8.
=20%-8.372%=11.628%